# Percolation on graphs of polynomial growth is local: analyticity, supercritical sharpness, isoperimetry

### v1

@@ 2.4 Cutconnectivity and quasi-isometries
Let us consider some $N>0$ that will be chosen along the way to be large enough depending on $p$. We will estimate $\mathbb{P}_{p}(|\mathcal{C}_{x}^{\mathrm{bad}}(N)|\geqslant n)$ using a union bound. To this end, let $\mathscr{I}$ be a $t$-connected set of vertices that contains $x$ such that $|\mathscr{I}|=n$. Write $r=10N+1$. We say that a set of vertices $A$ is $2r$-separated if $d_{\mathscr{G}}(u,v)>2r$ for all distinct $u,v\in A$. Let $\mathscr{I}^{\prime}\subset\mathscr{I}$ be a $2r$-separated set that is maximal for inclusion among $2r$-separated subsets of $\mathscr{I}$. Notice that by maximality, for every $y\in\mathscr{I}$, there exists $z\in\mathscr{I}^{\prime}$ such that $y\in B_{2r}(z)$. Hence, $|\mathscr{I}^{\prime}|\geqslant|\mathscr{I}|/|B_{2r}|$.
We claim that $\mathscr{I}^{\prime}$ is $(4r+t)$-connected. Indeed, consider two vertices $u,v\in\mathscr{I}^{\prime}$ and let $z,w\in\mathscr{I}$ be two vertices such that $d_{\mathscr{G}}(u,z)\leqslant 2r$ and $d_{\mathscr{G}}(v,w)\leqslant 2r$. As $\mathscr{I}$ is $t$-connected, we can pick a $t$-connected path $\gamma$ in $\mathscr{I}$ from $z$ to $w$, i.e. consecutive vertices in $\gamma$ are at distance at most $t$ apart in $\mathscr{G}$. Consider two consecutive vertices $a,b$ in $\gamma$, and pick two vertices $a^{\prime},b^{\prime}\in\mathscr{I}^{\prime}$ such that $d_{\mathscr{G}}(a,a^{\prime})\leqslant 2r$ and $d_{\mathscr{G}}(b,b^{\prime})\leqslant 2r$. Such vertices exist by the maximality of $\mathscr{I}^{\prime}$. We then have $d_{\mathscr{G}}(a^{\prime},b^{\prime})\leqslant 4r+t$. Iterating this procedure we can construct a $(4r+t)$-connected path from $u$ to $v$. This proves the claim.
We can thus conclude that when the event $\{|\mathcal{C}_{x}^{\mathrm{bad}}(N)|\geqslant n\}$ happens, there exists a $(4r+t)$-connected set of vertices $\mathcal{S}$ that is $2r$-separated, has cardinality $|\mathcal{S}|\geqslant n/|B_{2r}|$, contains a vertex in $B_{2r}(x)$, and such that every vertex in $\mathcal{S}$ is $N$-bad. The union bound gives
@@
Recall that either $\mathscr{I}$ contains $o$ or $o$ lies in a finite component of its complement in $\mathscr{G}$. This implies that $\mathscr{I}$ intersects both infinite geodesic rays $(\gamma(0),\gamma(1),\ldots)$ and $(\ldots,\gamma(-1),\gamma(0))$ at some vertices $x^{+}$ and $x^{-}$, respectively. Since $\mathscr{I}$ is $t$-connected and $|\mathscr{I}|=n$, it follows that $d_{\mathscr{G}}(x^{+},x^{-})\leqslant tn$, which implies that $x^{+}=\gamma(i)$ for some $0\leqslant i\leqslant t(n-1)$. ∎
@@
Let $r=10N+1$, and recall that $\mathscr{H}^{(i)}$ is the graph with vertex set $V$, where we draw an edge between $u$ and $v$ if $1\leqslant d_{\mathscr{H}}(u,v)\leqslant i$. Recall also that a set of vertices $A$ is $2r$-separated if $d_{\mathscr{G}}(u,v)>2r$ for all distinct $u,v\in A$. Given a multi-interface $\mathscr{M}=\{\mathscr{I}_{1},\mathscr{I}_{2},\ldots,\mathscr{I}_{k}\}$ of size $n$, let $S_{1},S_{2},\ldots,S_{\ell}$ be the connected components of $\bigcup_{j=1}^{k}\mathscr{I}_{j}$ in the graph $\mathscr{H}^{(2r)}$, i.e. each $S_{j}$ is $2r$-connected, and is maximal for inclusion with respect to this property. Now, for each $S_{j}$, let $S^{\prime}_{j}\subset S_{j}$ be a $2r$-separated set that is maximal for inclusion with respect to this property. Then $S^{\prime}:=\bigcup_{j=1}^{\ell}S^{\prime}_{j}$ is $2r$-separated, and
@@
By arguing as in the proof of Proposition 3.6, we see that each $S^{\prime}_{j}$ is $6r$-connected.
@@ 4.2 Probability that a cluster is large and finite
which implies that $\delta\geqslant\varepsilon_{0}/2$. Since $\partial B_{k}$ is $t$-connected, we can now open all edges in the $t$-neighbourhood of $\partial B_{k}$ to connect all clusters intersecting $\partial B_{k}$. By taking the intersection with the event $\{o\longleftrightarrow\partial B_{k}\}$ and applying the Harris–FKG inequality we obtain that
@@
One of the biggest open problems in percolation theory is whether there is an infinite cluster at criticality for Bernoulli percolation on $\mathbb{Z}^{d}$ ($d\geqslant 2$) with the usual nearest neighbour structure. This question has been answered in the negative in dimension $2$ [Kes80] and in dimensions $d\geqslant 11$ [HS90, FvdH17], but remains open in the intermediate dimensions. Recently, this question has attracted interest beyond the setting of the hypercubic lattice.
Recall that we define $\mathfrak{A}$ to be the set of all graphs that can be written as a product of finitely many cycle-graphs with at least two infinite factors; see page 1.8. Let us call such a graph a slab if the number of infinite factors is exactly 2. In [DCST16], it is proved that slabs satisfy $\theta(p_{c})=0$. See also [BLPS99, Tim06, PPS06, Hut16, HH21] for related results on graphs growing faster than $\mathbb{Z}^{d}$.
@@
Let $\mathscr{G}\in\mathfrak{A}$. If $\mathscr{G}$ is the product of the square lattice with finitely many finite cycles (possibly none), then it follows from [DCST16] that $\theta_{\mathscr{G}}$ is continuous. Otherwise, $\mathscr{G}$ can be written as $\mathbb{Z}^{2}\times\mathbb{Z}^{k}\times\mathscr{H}$ with $k\geqslant 1$ and $\mathscr{H}$ a finite product of finite cycles. In this case, we can introduce the graph sequence $\mathscr{G}_{n}$ corresponding to $\mathbb{Z}^{2}\times(\mathbb{Z}/n\mathbb{Z})^{k}\times\mathscr{H}$, which indeed satisfies $\mathscr{G}_{n}\xrightarrow[n\to\infty]{}\mathscr{G}$.
@@
  • the article [DCST16] does not deal with all two-dimensional transitive graphs but only such graphs having all symmetries of the square lattice.

### v2

@@ 2.4 Cutconnectivity and quasi-isometries
Let us consider some $N>0$ that will be chosen along the way to be large enough depending on $p$. We will estimate $\mathbb{P}_{p}(|\mathcal{C}_{x}^{\mathrm{bad}}(N)|\geqslant n)$ using a union bound. To this end, let $\mathscr{I}$ be a $t$-coarse connected set of vertices that contains $x$ such that $|\mathscr{I}|=n$. Write $r=10N+1$. We say that a set of vertices $A$ is $2r$-separated if $d_{\mathscr{G}}(u,v)>2r$ for all distinct $u,v\in A$. Let $\mathscr{I}^{\prime}\subset\mathscr{I}$ be a $2r$-separated set that is maximal for inclusion among $2r$-separated subsets of $\mathscr{I}$. Notice that by maximality, for every $y\in\mathscr{I}$, there exists $z\in\mathscr{I}^{\prime}$ such that $y\in B_{2r}(z)$. Hence, $|\mathscr{I}^{\prime}|\geqslant|\mathscr{I}|/|B_{2r}|$.
We claim that $\mathscr{I}^{\prime}$ is $(4r+t)$-coarse connected. Indeed, consider two vertices $u,v\in\mathscr{I}^{\prime}$ and let $z,w\in\mathscr{I}$ be two vertices such that $d_{\mathscr{G}}(u,z)\leqslant 2r$ and $d_{\mathscr{G}}(v,w)\leqslant 2r$. As $\mathscr{I}$ is $t$-coarse connected, we can pick a $t$-coarse connected path $\gamma$ in $\mathscr{I}$ from $z$ to $w$, i.e. consecutive vertices in $\gamma$ are at distance at most $t$ apart in $\mathscr{G}$. Consider two consecutive vertices $a,b$ in $\gamma$, and pick two vertices $a^{\prime},b^{\prime}\in\mathscr{I}^{\prime}$ such that $d_{\mathscr{G}}(a,a^{\prime})\leqslant 2r$ and $d_{\mathscr{G}}(b,b^{\prime})\leqslant 2r$. Such vertices exist by the maximality of $\mathscr{I}^{\prime}$. We then have $d_{\mathscr{G}}(a^{\prime},b^{\prime})\leqslant 4r+t$. Iterating this procedure we can construct a $(4r+t)$-coarse connected path from $u$ to $v$. This proves the claim.
We can thus conclude that when the event $\{|\mathcal{C}_{x}^{\mathrm{bad}}(N)|\geqslant n\}$ happens, there exists a $(4r+t)$-coarse connected set of vertices $\mathcal{S}$ that is $2r$-separated, has cardinality $|\mathcal{S}|\geqslant n/|B_{2r}|$, contains a vertex in $B_{2r}(x)$, and such that every vertex in $\mathcal{S}$ is $N$-bad. The union bound gives
@@
Recall that either $\mathscr{I}$ contains $o$ or $o$ lies in a finite component of its complement in $\mathscr{G}$. This implies that $\mathscr{I}$ intersects both infinite geodesic rays $(\gamma(0),\gamma(1),\ldots)$ and $(\ldots,\gamma(-1),\gamma(0))$ at some vertices $x^{+}$ and $x^{-}$, respectively. Since $\mathscr{I}$ is $t$-coarse connected and $|\mathscr{I}|=n$, it follows that $d_{\mathscr{G}}(x^{+},x^{-})\leqslant tn$, which implies that $x^{+}=\gamma(i)$ for some $0\leqslant i\leqslant t(n-1)$. ∎
@@
Let $r=10N+1$, and recall that $\mathscr{H}^{(i)}$ is the graph with vertex set $V$, where we draw an edge between $u$ and $v$ if $1\leqslant d_{\mathscr{H}}(u,v)\leqslant i$. Recall also that a set of vertices $A$ is $2r$-separated if $d_{\mathscr{G}}(u,v)>2r$ for all distinct $u,v\in A$. Given a multi-interface $\mathscr{M}=\{\mathscr{I}_{1},\mathscr{I}_{2},\ldots,\mathscr{I}_{k}\}$ of size $n$, let $S_{1},S_{2},\ldots,S_{\ell}$ be the connected components of $\bigcup_{j=1}^{k}\mathscr{I}_{j}$ in the graph $\mathscr{H}^{(2r)}$, i.e. each $S_{j}$ is $2r$-coarse connected, and is maximal for inclusion with respect to this property. Now, for each $S_{j}$, let $S^{\prime}_{j}\subset S_{j}$ be a $2r$-separated set that is maximal for inclusion with respect to this property. Then $S^{\prime}:=\bigcup_{j=1}^{\ell}S^{\prime}_{j}$ is $2r$-separated, and
@@
By arguing as in the proof of Proposition 3.6, we see that each $S^{\prime}_{j}$ is $6r$-coarse connected.
@@ 4.2 Probability that a cluster is large and finite
which implies that $\delta\geqslant\varepsilon_{0}/2$. Since $\partial B_{k}$ is $t$-coarse connected, we can now open all edges in the $t$-neighbourhood of $\partial B_{k}$ to connect all clusters intersecting $\partial B_{k}$. By taking the intersection with the event $\{o\longleftrightarrow\partial B_{k}\}$ and applying the Harris–FKG inequality we obtain that
@@
4.3 Exponential cluster repulsion and isoperimetry
In this section, our aim is to prove that the probability $\mathcal{C}_{o}$ is finite but touches the infinite cluster at $t$ places decays exponentially in $t$. As a corollary, we obtain an isoperimetric inequality for the infinite cluster.
Let us first give the formal definition. Consider a graph $\mathscr{G}\in\mathfrak{G}$. For two distinct clusters $\mathcal{C}$ and $\mathcal{C}^{\prime}$, a touching edge is an edge of $\mathscr{G}$ with one endpoint in $\mathcal{C}$ and the other in $\mathcal{C}^{\prime}$. The number of such edges is denoted $\tau(\mathcal{C},\mathcal{C}^{\prime})$. For supercritical percolation on $\mathscr{G}$, the almost surely unique infinite cluster is denoted $\mathcal{C}_{\infty}$.
We now state and prove the following result.
Theorem 4.5.
$\forall p\in[p_{0},1],\quad\mathbb{P}_{p,\mathscr{H}}\big(|\mathcal{C}_{o}|<\infty,\penalty 10000\ \tau(\mathcal{C}_{o},\mathcal{C}_{\infty})\geqslant k\big)\leqslant\exp(-ck).$
Proof.
Let $p_{0}\in(p_{c}(\mathscr{G}),1]$. Arguing as in the proof of Theorem 1.2, we see that it suffices to find an $\varepsilon>0$ such that the desired uniform exponential decay holds for every $p\in(p_{0}-\varepsilon,p_{0}+\varepsilon)\cap[0,1]$.
To this end, consider some large enough constants $N,r>0$ and a graph $\mathscr{H}\in\mathfrak{G}$ such that $R(\mathscr{G},\mathscr{H})\geqslant r$. Consider a configuration $\omega$ such that $\mathcal{C}_{o}$ is finite, and note that $|\mathcal{C}_{o}|\geqslant\tau(\mathcal{C}_{o},\mathcal{C}_{\infty})/D$, where $D$ is degree of $\mathscr{G}$. Hence, if $\tau(\mathcal{C}_{o},\mathcal{C}_{\infty})>D|B_{20N,\mathscr{G}}|$ and $r\geqslant 20N$, then $\textup{diam}(\mathcal{C}_{o})>20N$. Consider the associated interface $\mathcal{I}$ and recall that the closed edges of $\overline{\mathcal{I}}$ form a bond-cutset between $o$ and infinity by Lemma 3.3. Let us denote this bond-cutset $\Pi$. Then $\Pi$ needs to contain all touching edges between $\mathcal{C}_{o}$ and $\mathcal{C}_{\infty}$. Hence we have
$|\mathcal{I}|\geqslant\frac{|\overline{\mathcal{I}}|}{|B_{10N,\mathscr{G}}|}\geqslant\frac{\tau(\mathcal{C}_{o},\mathcal{C}_{\infty})}{D|B_{10N,\mathscr{G}}|}.$
The desired result follows from Proposition 3.8 for large values of $k$. Finally, since we have $|\mathcal{C}_{o}|\geqslant\tau(\mathcal{C}_{o},\mathcal{C}_{\infty})/D$, we can use Theorem 4.2 to handle the small values of $k$. ∎
Given a finite subset $S$ of $\mathcal{C}_{\infty}$, we distinguish between two notions of edge boundary in $\mathcal{C}_{\infty}$. Following Pete’s notation [PET08], we denote $\partial^{+}_{\mathcal{C}_{\infty}}S$ the set of open edges of $\mathcal{C}_{\infty}$ with one endpoint in $S$ and the other in the infinite connected component of $\mathcal{C}_{\infty}\setminus S$. We also denote $\tilde{\partial}_{\mathcal{C}_{\infty}}S$ the set of edges in $\mathscr{G}$ with one endpoint in $S$ and the other in the infinite connected component of $\mathcal{C}_{\infty}\setminus S$.
We now prove the following anchored isoperimetric inequality for the infinite cluster. The proof is similar to that of [PET08, Theorem 1.2].
Theorem 4.6.
Let $\mathscr{G}\in\mathfrak{G}$ and $p_{0}\in(p_{c}(\mathscr{G}),1]$. Then, there is a constant $c>0$ such that the set of all $\mathscr{H}\in\mathfrak{G}$ satisfying $p_{c}(\mathscr{H})<p_{0}$ and the following condition is a neighbourhood of the graph $\mathscr{G}$: for every $p\in[p_{0},1]$, there exists $\alpha>0$ such that for all $M\geqslant 1$, we have
$\mathbb{P}_{p}\left(\exists S\text{ connected}:o\in S\subset\mathcal{C}_{\infty},\ M\leqslant|S|<\infty,\ \frac{|\partial^{+}_{\mathcal{C}_{\infty}}S|}{|S|^{1-1/d(\mathscr{H})}}\leqslant\alpha\right)\leqslant\exp\left(-cM^{1-1/d(\mathscr{H})}\right).$
Remark 4.7.
We note that the value of $\alpha$ that the argument gives depends on $p$ and degenerates to 0 as $p$ tends to $1$, while the value of $c$ depends only on $p_{0}$.
Proof.
The result holds for $p=1$ by Proposition 4.3, so let us assume that $p_{0}\in(p_{c}(\mathscr{G}),1)$ and prove the result for $p\in[p_{0},1)$. One can handle the case where $|S|=m$ for some small $m$ by opening all edges in the ball of radius $2m$ and forcing the ball to connect to infinity to ensure that $\frac{|\partial^{+}_{\mathcal{C}_{\infty}}S|}{|S|^{1-1/d(\mathscr{H})}}$ remains bounded away from 0, so it remains to handle the case where $M$ is large enough.
Define the event
$\mathcal{Y}(m,t):=\left\{|\mathcal{C}_{o}|=m\quad\text{and}\quad\tau(\mathcal{C}_{o},\mathcal{C}_{\infty})=t\right\}.$
By Theorems 4.2 and 4.5, there exist $r,c>0$ such that for every graph $\mathscr{H}\in\mathfrak{G}$ satisfying $R(\mathscr{G},\mathscr{H})\geqslant r$ and every $p\geqslant p_{0}$, we have
$\mathbb{P}_{p}(\mathcal{Y}(m,t))\leqslant\exp\bigl(-c\max\{m^{1-1/d},t\}\bigr),$
where $d$ denotes the dimension of $\mathscr{H}$. We now introduce the events
$\mathcal{X}(m,t,s):=\left\{\exists S\text{ connected}:o\in S\subset\mathcal{C}_{\infty},\ |S|=m,\ |\tilde{\partial}_{\mathcal{C}_{\infty}}S|=t,\ |\partial^{+}_{\mathcal{C}_{\infty}}S|=s\right\}.$
and
$\mathcal{E}_{\alpha,M}:=\left\{\exists S\text{ connected}:o\in S\subset\mathcal{C}_{\infty},\ M\leqslant|S|<\infty,\ \frac{|\partial^{+}_{\mathcal{C}_{\infty}}S|}{|S|^{1-1/d}}\leqslant\alpha\right\}.$
By a union bound, and ignoring integer parts to simplify the notation, we have
$\mathbb{P}_{p}\left(\mathcal{E}_{\alpha,M}\right)\leqslant\sum_{m\geqslant M}\sum_{t=1}^{Dm}\sum_{s=1}^{c_{m,t}}\mathbb{P}_{p}(\mathcal{X}(m,t,s)),$ (13)
where $c_{m,t}=\min\{\alpha m^{1-1/d},t\}$, and $d,D$ are the dimension and degree of $\mathscr{H}$, respectively. We now bound $\mathbb{P}_{p}(\mathcal{X}(m,t,s))$ using the multi-valued map principle.
Given a configuration $\xi\in\{0,1\}^{E(B_{2m})}$, we let $\mathcal{X}(m,t,s,\xi)$ denote the event $\{\omega|_{E(B_{2m})}=\xi\text{ and }\omega\in\mathcal{X}(m,t,s)\}$. For a $\xi$ such that $\mathbb{P}_{p}(\mathcal{X}(m,t,s,\xi))>0$ and a finite connected set $S\ni o$, define a new event $F(\xi,S)\subset\mathcal{Y}(m,t)$ by closing the edges in $\partial^{+}_{\mathcal{C}_{\infty}}S$. Note that $\mathbb{P}(F(\xi,S))=Q^{s}\mathbb{P}(\mathcal{X}(m,t,s,\xi))$, where $Q=\frac{p}{1-p}$. For each $\omega^{\prime}$ in the image of $F$ such that $\omega^{\prime}\in\mathcal{Y}(m,t)$, there are $t\choose s$} pre-images $(\xi,S)$ under $F$. Hence, we have
$\mathbb{P}_{p}(\mathcal{X}(m,t,s))\leqslant{t\choose s}Q^{s}\mathbb{P}_{p}(\mathcal{Y}(m,t)).$
With these bounds at hand, we estimate $\sum_{t=1}^{Dm}\sum_{s=1}^{c_{m,t}}\mathbb{P}_{p}(\mathcal{X}(m,t,s))$ by splitting it into $2$ sums according to whether $t\leqslant K\alpha m^{1-1/d}$ or $t>K\alpha m^{1-1/d}$, where $K=Q+2+4/c$. To this end, note that for $\alpha>0$ small enough to satisfy $K\alpha\log(1+Q)<c/2$, we have that for every $m$ large enough,
$\sum_{t=1}^{K\alpha m^{1-1/d}}\;\sum_{s=1}^{c_{m,t}}\ \mathbb{P}_{p}(\mathcal{X}(m,t,s))\leqslant\sum_{t=1}^{K\alpha m^{1-1/d}}(1+Q)^{t}\exp\bigl(-cm^{1-1/d}\bigr)\leqslant\exp\left(-\frac{c}{2}m^{1-1/d}\right).$ (14)
For the second sum, note that for $t>K\alpha m^{1-1/d}>2\alpha m^{1-1/d}$ and $s\leqslant\alpha m^{1-1/d}$, we have
${t\choose s}Q^{s}\leqslant{t\choose\alpha m^{1-1/d}}(1+Q)^{s}\leqslant\left(\frac{et}{\alpha m^{1-1/d}}\right)^{\alpha m^{1-1/d}}(1+Q)^{\alpha m^{1-1/d}},$
hence
$\sum_{s=1}^{\alpha m^{1-1/d}}{t\choose s}Q^{s}\leqslant\alpha m^{1-1/d}\exp\Bigl(\alpha\bigl(1+\log(\beta/\alpha)+\log(1+Q)\bigr)\,m^{1-1/d}\Bigr),$
where $\beta=t/m^{1-1/d}$. Since $\beta/\alpha>1+Q$ and $\alpha<cK\alpha/4<c\beta/4$, if $\alpha$ is sufficiently small, then we have
$\alpha(1+\log(\beta/\alpha)+\log(1+Q))<\alpha(1+2\log(\beta/\alpha))<c\beta/2.$
Hence, for every $m$ large enough, we get
$\sum_{t=1+Q\alpha m^{1-1/d}}^{Dm}\;\sum_{s=1}^{c_{m,t}}\ \mathbb{P}_{p}(\mathcal{X}(m,t,s))\leqslant\sum_{t>K\alpha m^{1-1/d}}\exp\left(-\frac{c}{2}t\right).$ (15)
The desired result follows from putting together (13), (14) and (15). ∎
Remark 4.8.
The above theorem implies in particular the following result: for every $\mathscr{G}\in\mathfrak{G}$, the anchored isoperimetric dimension of the cluster of the origin is equal to the dimension of $\mathscr{G}$ as soon as this cluster is infinite. This result for fixed graphs can alternatively be obtained from arguments in [PET08] and [CMT24]. The parallel work [AK26] provides a more general result covering long-range percolation for fixed graphs in $\mathfrak{G}$ for which supercritical sharpness holds; see their Theorem 1.5. Still, we believe that the viewpoint of interfaces adopted in the present paper is a relevant way to revisit the techniques from [PET08], even for the hypercubic lattice $\mathbb{Z}^{d}$.
@@ 4.3 Exponential cluster repulsion and isoperimetry
One of the biggest open problems in percolation theory is whether there is an infinite cluster at criticality for Bernoulli percolation on $\mathbb{Z}^{d}$ ($d\geqslant 2$) with the usual nearest neighbour structure. This question has been answered in the negative in dimension $2$ [KES80] and in dimensions $d\geqslant 11$ [HS90, Fv17], but remains open in the intermediate dimensions. Recently, this question has attracted interest beyond the setting of the hypercubic lattice.
Recall that we define $\mathfrak{A}$ to be the set of all graphs that can be written as a product of finitely many cycle-graphs with at least two infinite factors; see page 1.9. Let us call such a graph a slab if the number of infinite factors is exactly 2. In [DST16], it is proved that slabs satisfy $\theta(p_{c})=0$. See also [BLP+99, TIM06, PPS06, HUT16, HH21] for related results on graphs growing faster than $\mathbb{Z}^{d}$.
@@
Let $\mathscr{G}\in\mathfrak{A}$. If $\mathscr{G}$ is the product of the square lattice with finitely many finite cycles (possibly none), then it follows from [DST16] that $\theta_{\mathscr{G}}$ is continuous. Otherwise, $\mathscr{G}$ can be written as $\mathbb{Z}^{2}\times\mathbb{Z}^{k}\times\mathscr{H}$ with $k\geqslant 1$ and $\mathscr{H}$ a finite product of finite cycles. In this case, we can introduce the graph sequence $\mathscr{G}_{n}$ corresponding to $\mathbb{Z}^{2}\times(\mathbb{Z}/n\mathbb{Z})^{k}\times\mathscr{H}$, which indeed satisfies $\mathscr{G}_{n}\xrightarrow[n\to\infty]{}\mathscr{G}$.
@@
  • the article [DST16] does not deal with all two-dimensional transitive graphs but only such graphs having all symmetries of the square lattice.